OpenStudy (anonymous):

Use the graph to find the solution to the system of equations. Enter your answer as an ordered pair. y = -x + 3 y = x - 3

OpenStudy (anonymous):

-x + 3=x - 3 find x value net put it in 1st equation 2 get y val

OpenStudy (anonymous):

@sail Do i subtract or add 3 to the left or right side of the equation?

OpenStudy (anonymous):

@sai1234

OpenStudy (anonymous):

add 3 on both sides

OpenStudy (anonymous):

@sai1234 so the equation will look like -x+6=0 ? after i add 3 on both sides?

OpenStudy (anonymous):

@sai1234 so x=-6 ?